OpenStudy (anonymous):

Find the derivatives of the function f(x) = 4sqrtx + 3/x^4− 8x^5/2 + 3x − 4

OpenStudy (anonymous):

This function f is a sum 5 terms. To compute the derivative of f you need to compute the derivative of each of these terms add add them together.

OpenStudy (anonymous):

So do you know the derivative of the first term 4sqrt(x)?

OpenStudy (anonymous):

no not really I just start learning it today

OpenStudy (anonymous):

Ok, well the first thing you need to know is the derivative of a power of: \[\frac{d}{dx}x^c=c x^{c-1} \] d/dx is notation for the derivative.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you apply that to the fourth term, 3x?

OpenStudy (anonymous):

maybe I should've put it like this: \[\frac{d}{dx} a x^c=acx^{c-1}\] for a constant a.

OpenStudy (anonymous):

You're having a hard time finding the derivative of 3x?

OpenStudy (anonymous):

ok see what confuses me my professor haven't even taught this

OpenStudy (anonymous):

What did your professor teach you?

OpenStudy (anonymous):

never mind I got the problem now. we didn't have to do this one my bad.

OpenStudy (anonymous):

all right then